OpenStudy (anonymous):

Find the root and vertex of this parabola. x^2 + 7X + 10 = 0 thanks

OpenStudy (anonymous):

i don't know about the vertex but, that formula can be factored out into (x+2)(x+5)

OpenStudy (anonymous):

so your roots would be -2, -5

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

vertex can be found using f(-b/2a) = x^2 + 7X + 10 ax^2 + bx + C = 0 Where b = 7 and a = 1

OpenStudy (amistre64):

also the vertex axis is midpoint between roots for a parabola